Sang invested some money at 13% interest. Sang also invested $66 more than 3 times that amount at 12%. How much is invested at each rate if Sang receives $529.28 in interest after one year?
Answer by ikleyn(52802)

.
Sang invested some money at 13% interest. Sang also invested $66 more than 3 times that amount at 12%. How much is invested at each rate if Sang receives $529.28 in interest after one year?
~~~~~~~~~~~~~~~~~~~~~~~~~~

0.13x + 0.12y = 529.28    (1)   (x = an amount at 13%,
y = 3x + 66               (2)    y = an amount at 12%).

if you want to have the system in two unknowns, or

0.13x + 0.12(3x+66) = 529.28,   (3)

if you want to have the system in one single unknown.

To solve (3), first multiply (3) by 100 (both sides). You will have

13x + 12(3x+66) = 52928.

Solve it for x.